# Scaling Laws of Synthetic Data for Language Models

### v2

@@ Appendix B Examples about Extracted Topics and Key Concepts from Reference Document
<level>College</level> <subject>Vector Calculus</subject> <topic> Topics: 1. Surface Integrals of Vector Fields 2. Oriented Surfaces 3. Unit Normal Vectors 4. Flux of a Vector Field 5. Parametric Surfaces </topic> <key_concept> Key Concepts: 1. Surface Integrals of Vector Fields:       1.1. Definition of surface integral of a vector field       1.2. Flux of a vector field across a surface       1.3. Application of surface integrals in fluid dynamics       1.4. Evaluation of surface integrals using parametric surfaces       1.5. Surface integrals over closed surfaces 2. Oriented Surfaces:       2.1. Definition of an oriented surface       2.2. Positive and negative orientations       2.3. Unit normal vectors and their role in orientation       2.4. Orientation conventions for closed surfaces       2.5. Impact of orientation on surface integrals 3. Unit Normal Vectors:       3.1. Definition and calculation of unit normal vectors       3.2. Gradient vector and its role in finding normal vectors       3.3. Normal vectors for surfaces given by z = f(x, y)       3.4. Normal vectors for parametric surfaces       3.5. Adjusting normal vectors to match desired orientation 4. Flux of a Vector Field:       4.1. Definition of flux       4.2. Physical interpretation of flux in fluid dynamics       4.3. Calculation of flux using surface integrals       4.4. Flux across closed surfaces       4.5. Application of flux in Gauss’s Law 5. Parametric Surfaces:       5.1. Definition and representation of parametric surfaces       5.2. Calculation of normal vectors for parametric surfaces       5.3. Evaluation of surface integrals using parametric surfaces       5.4. Parameterization of common surfaces (e.g., spheres, cylinders)       5.5. Conversion between parametric and non-parametric forms </key_concept>

### v3

@@ Appendix E Examples about Extracted Topics and Key Concepts from Reference Document
<level>College</level> <subject>Vector Calculus</subject> <topic> Topics: 1. Surface Integrals of Vector Fields 2. Oriented Surfaces 3. Unit Normal Vectors 4. Flux of a Vector Field 5. Parametric Surfaces </topic> <key_concept> Key Concepts: 1. Surface Integrals of Vector Fields:    $1.$ Definition of surface integral of a vector field    1.2. Flux of a vector field across a surface    1.3. Application of surface integrals in fluid dynamics    1.4. Evaluation of surface integrals using parametric surfaces    1.5. Surface integrals over closed surfaces 2. Oriented Surfaces:    2.1. Definition of an oriented surface    $2.$ Positive and negative orientations    2.3. Unit normal vectors and their role in orientation    2.4. Orientation conventions for closed surfaces    2.5. Impact of orientation on surface integrals 3. Unit Normal Vectors:    3.1. Definition and calculation of unit normal vectors    3.2. Gradient vector and its role in finding normal vectors    $3.$ Normal vectors for surfaces given by z = f(x, y)    3.4. Normal vectors for parametric surfaces    3.5. Adjusting normal vectors to match desired orientation 4. Flux of a Vector Field:    4.1. Definition of flux    4.2. Physical interpretation of flux in fluid dynamics    4.3. Calculation of flux using surface integrals    $4.$ Flux across closed surfaces    4.5. Application of flux in Gauss’s Law 5. Parametric Surfaces:    5.1. Definition and representation of parametric surfaces    5.2. Calculation of normal vectors for parametric surfaces    5.3. Evaluation of surface integrals using parametric surfaces    5.4. Parameterization of common surfaces (e.g., spheres, cylinders)    $5.$ Conversion between parametric and non-parametric forms </key_concept>